IQ scores of people around the World are normally distributed, with a mean of 100 and a standard deviation of 15. A genius is someone with an IQ greater than or equal to 140. What percent of the population is considered genius ?
1
Expert's answer
2020-06-01T18:26:00-0400

Since IQ of people has a normal distribution with "\mu=100,\,\,\sigma=15" , its probability density function (see e.g., https://en.wikipedia.org/wiki/Normal_distribution) is

"p(x)=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac12(\frac{x-\mu}{\sigma})^2}" , where "\mu=100,\,\,\sigma=15" .

We compute the following integral:

"\int_{140}^{+\infty}p(x)dx=\frac{1}{15\sqrt{2\pi}}\int_{140}^{+\infty}e^{-\frac12(\frac{x-100}{15})^2}dx\approx0.0038"

Then, the percent of people with IQ greater than or equal to 140 is approximately 0.0038*100%=0.38%

Answer:0.38%


P.S. the integral was computed with a help of Anaconda (https://www.anaconda.com/products/individual).

A code for the computation is:


from scipy import integrate

import numpy as np

import math

func = lambda x:(1/(15*math.sqrt(2)*math.sqrt(math.pi)))*math.exp(-1/2*((x-100)/15)*

((x-100)/15))


Pr = integrate.quad(func, 140, np.Infinity)

print(Pr)